A can do a job in 6 days while B can do a job in 8 days. They worked together for 2 days when B left the job. How much time will A require to complete the work?
Answer by richwmiller(17219)

2/6+2/8+x/6=1
1/3+1/4+x/6=1
x/6+7/12 = 1
x/6=5/12
2x/12=5/12
2x=5
x=2.5 more hours to finish